a 6m ladder leans against a house so thats its base is 2 metres out from the bottom of the house how far up the house does the ladder reach
Answer by checkley71(8403)

2^2+X^2=6^2
4+X^2=36
X^2=36-4
X^2=32
X=SQRT32
X=5.65685M IS THE HEIGHT OF THE LADDER TIOUCHING THE HOUSE.